OpenStudy (anonymous):

*screams with frustration* Ok so irritated. I really can't get this ^3sqrt(135x^8)...simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers.

OpenStudy (saifoo.khan):

\[\large \sqrt[3]{3 \times 3 \times 3 \times 5 x^8}\]

OpenStudy (saifoo.khan):

Can u solve now?

OpenStudy (anonymous):

so multiply whats in the radical?

OpenStudy (saifoo.khan):

No, i factored out 135. Or in other words, broke it into pieces so we can simplify it.

OpenStudy (saifoo.khan):

Now since it's a cuteroot, "3" comes out, leaving behind, \[\large 3 \sqrt[3]{5 \times x^8}\]

OpenStudy (anonymous):

Oh ok I see what you did now. thanks for explaining that for me.

OpenStudy (saifoo.khan):

Finally, solve the variable, we will get. \[\Large 3 x^2 \sqrt[3]{5 x^2}\]

OpenStudy (saifoo.khan):

Welcome. Please hit good answer. Thanks.

OpenStudy (anonymous):

One more question if anyone can answer it...on the answer saifoo gave am I putting a multiplication symble between the 3x^2 and radical?

OpenStudy (saifoo.khan):

No, that's "x" the variable.

OpenStudy (saifoo.khan):

Multiplications signs aren't needed if you are doing this online.

OpenStudy (anonymous):

Ok thank you.